Andrew is paid £250 a week.
Each week, he
shares his pay with his sister in the ratio 3:2
saves 12% of his share.
How many weeks will it take Andrew to save £360?

Answer by Theo(13342)

he gets paid 250 a week.
he shares his pay with his sister in the ratio of 3:2.
presumably he keeps the 3 and gives her the 2.
take 250 and divide it by 3 + 2 = 5 and you get 50.
he keeps 3 * 50 = 150 and he gives her 2 * 50 = 100.
he saves 12% of 150 = 150 * .12 = 18 each week.
it will take him 360 / 18 = 20 weeks to save 360.
20 * 18 = 360.